OpenStudy (anonymous):

find the slope y= -x + 3

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

thank you

OpenStudy (agreene):

y=mx+b y=-x+3 or y=(-1)x+3 m=-1

OpenStudy (anonymous):

yw